OpenStudy (anonymous):

g(x) = √5x - 4 h(x) = 4x2 + 7 Find: h(g(1))

OpenStudy (anonymous):

is the g(x) \[\sqrt{5x-4} or \sqrt{5}x-4\]

OpenStudy (anonymous):

So let's assume you have g(x) = \[\sqrt{5x-4}\] let's find g(1) = \[\sqrt{5*1-4}\] = \[\sqrt{1}\] = 1 g(1) = 1 so, h(1) = 4*1*2+7 = 56 you have mistakes in your question by the way, what do you mean by 4x2? I solved it as if you had something like 8x+7

OpenStudy (precal):